OpenStudy (anonymous):

[-(30-x)]^2 what is the answer to this, please show all steps.

OpenStudy (anonymous):

(30-x)^2 [-x]^2=x^2 so if x=(30-x) a:(30-x)^2

OpenStudy (e.cociuba):

For the result I got (x-30)^2. Want an explanation?

OpenStudy (anonymous):

would love an explanation.

OpenStudy (anonymous):

should it not be a quadratic in the end?

OpenStudy (anonymous):

expand it out.

OpenStudy (e.cociuba):

Yes when you put it in expanded form u get: